# Feeding the Crave: How People with Eating Disorders Get Trapped in the Perpetual Cycle of Digital Food Content

### v3

@@ Abstract.
Recent studies have examined how digital food content impacts viewers’ dietary health. A few have found that individuals with eating disorders are particularly sensitive to digital food content, such as eating and cooking videos, which contribute to disordered eating behaviors. However, there is a lack of comprehensive studies that investigate how these individuals interact with various digital food content. To fill this gap, we conducted two rounds of studies (N=23 and 22, respectively) with individuals with eating disorders to understand their motivations and practices of consuming digital food content. Our study reveals that participants anticipate positive effects from food media to overcome their condition, but in practice, it often exacerbates their disorder. We also discovered that many participants experienced a cycle of quitting and returning to digital food content consumption. Based on these findings, we articulate design implications for digital food content and multimedia platforms to support vulnerable individuals.
copyright: none
@@ 1. Introduction
Digital media’s exponential growth has made food content ubiquitous and diversified. On social media, the popularity of food videos (e.g., eating and cooking broadcasts) and images has significantly increased (Kang et al., 2020; Mejova et al., 2016; De Solier, 2005; Adema, 2000). On TikTok, the hashtag #TikTokFood has been viewed over 43 billion times (Kitchens, 2021), and Instagram has over 310 million posts tagged with #foodporn (Instagram, 2024). The online food market, especially food delivery services often showcasing visually enticing images, has also seen tremendous growth, more than threefold, since 2017, rising as a global market worth more than \$150 billion (Insights, 2023; Inc., 2021).
@@
A growing body of research has examined the impact of increased exposure to various digital food content on eating habits and dietary health (Petit et al., 2016; Kinard, 2016; Ofli et al., 2017; Vermeir and Roose, 2020; Pope et al., 2015; Choi et al., 2024; Wu et al., 2024). The trend of watching and posting food images and videos with catchy hashtags has been associated with inadequate eating habits (Yan and Corino, 2019; Qutteina et al., 2019), unhealthy food choices (Peng and JEMMOTT III, 2018; Basso et al., 2018), and overeating (Bodenlos and Wormuth, 2013; Halford et al., 2004) through various visual and social factors (Seal et al., 2022). Browsing and ordering food using online shopping and food delivery services is also known to involve unhealthy food choices (Yeo et al., 2017), unhealthy eating (Eu and Sameeha, 2021), and obesity (Stephens et al., 2020). Moreover, many digital food content features take-away food (Kang et al., 2020), and combining eating content with online food delivery services could lead to unhealthy lifestyles.
@@
Recent studies have highlighted that the negative impacts of digital food content on dietary health are more intense in individuals with eating disorders (EDs), who are particularly sensitive to such content (Jin, 2018; Legenbauer et al., 2004) and engage with it significantly more than the general population (Bachner-Melman et al., 2018). When coupled with frequent unconscious exposure to digital food content, digital food content could exacerbate ED symptoms, such as binge eating and obsessive food thoughts (Choi et al., 2024; Kinkel-Ram et al., 2022). Despite these findings, there remains a significant gap in understanding how individuals with EDs engage with digital food content and how such engagement ultimately contributes to their disordered eating in their daily lives. In this regard, we aim to shed light on the experience of people with disordered eating behaviors with digital food content by answering the following question: How do people with EDs perceive and interact with digital food content?
@@
To answer this question, we conducted two rounds of studies with individuals with eating disorders: one round of exploratory study (N=23) and another round of in-depth study (N=22). The exploratory study assesses how many individuals are concerned with associations of digital food content and disordered eating, and identifies which digital food content types are associated with disordered eating. Notably, we refrained from bringing up digital food content, allowing participants to broach the subject voluntarily. Sixteen participants out of twenty-three voluntarily linked digital food content to their ED symptoms. Our analysis identified two representative digital food content positively associated with ED symptoms: food media (e.g., images and videos) and food delivery apps.
To deepen our understanding of ED individuals’ motivation and practices of consuming each type of digital food content, we conducted in-depth interviews with 22 individuals (including two from our exploratory study) with EDs. Our main finding highlights a significant disparity between the expectations of individuals when using digital food content (e.g., vicarious satisfaction) and the actual outcomes they experience (e.g., a binge-eating stimulus). Amidst this disparity, many individuals have experienced a recurring cycle of quitting and returning to digital food content consumption behavior. In addition, we identified a unique subcategory of food media that some utilized as a self-care strategy: videos created by people with EDs (e.g., binge-eating broadcasts)
Our finding emphasizes the significant impact of digital food exposure on individuals with EDs, often resulting in developing or exacerbating disordered eating behaviors. The increasing presence of virtual food potentially triggering unhealthy eating habits fails to consider the growing occurrence of eating disorders, which is a concern for relevant organizations and policymakers (Taillie et al., 2019; Times, 2023; Rilloraza, 2024). Understanding the significance of visually and auditorily presented food in today’s context is crucial for guiding populations toward more appropriate food behaviors and choices. From our findings, we propose design implications for intervention to promote healthy interactions with digital food content.
@@
  • We identified two representative kinds of digital food content, food media and food delivery apps, associated with eating disorders in everyday life.
  • Through an in-depth study, we revealed the perceptions, motivations, impacts, and patterns of consuming digital food content among individuals with eating disorders.
  • Finally, we provided design implications to improve multimedia platforms for vulnerable populations and offered insights for future directions in digital food content curation.
  • Our study complements prior research on digital food content by focusing on individuals with eating disorders, providing unique insights into their digital content consumption patterns and implications for HCI designers.
@@ 2.2. An Overarching Presence of Digital Food Content
This omnipresence of digital food content underscores the need to explore its impact on eating behaviors, especially among individuals with eating disorders. The subsequent section delves into the intricate dynamics of digital food content’s influence on individuals’ eating behaviors, shedding light on its constructive and detrimental effects.
@@ 2.3. Double-sided Impact of Digital Food Content on Eating Behaviors
Digital food content exerts a complex influence on individuals’ eating behaviors, presenting a dual-sided impact characterized by both constructive and detrimental facets (Strand and Gustafsson, 2020). While many individuals engage with digital food content for inspiration, recipe ideas, and cultural connections to food (De Solier, 2005), the comprehensive impact of such media on individuals with eating disorders, who are much more sensitive to digital food content than the general population (Schienle et al., 2009; Lee et al., 2017), has received limited attention. The relationship between individuals’ consumption of digital food content, including the engagement with eating broadcasts (Mukbang), and the potential escalation of disordered eating behaviors has garnered recognition in prior research, primarily focusing on the general population (von Ash et al., 2023; Anari and Eghtesadi, 2023). However, these investigations have yielded varying insights and have yet to explicitly target individuals with eating disorders.
@@
Conversely, a growing body of research has highlighted the potentially detrimental impact of digital food content, especially when exacerbating unhealthy eating habits (Pope et al., 2015; Spence et al., 2016). The omnipresence of visually appealing, calorie-laden food imagery in digital content can escalate disordered eating behaviors (Bodenlos and Wormuth, 2013; Ventura et al., 2021). Studies have pointed out that exposure to images and videos of excessive food consumption, particularly those featuring eating challenges or binge eating, can inadvertently trigger or reinforce unhealthy eating patterns (Spence et al., 2019). Such content inadvertently glorifies or normalizes overconsumption (Kircaburun et al., 2021c). Additionally, the visual appeal of enticing food displays and the sensory experience of hearing eating sounds may inadvertently intensify cravings and compulsive eating tendencies (Spence et al., 2016; Anjani et al., 2020).
@@
The current research landscape lacks a profound understanding of how individuals with eating disorders engage with digital food content—a realm encompassing spontaneous interactions and passive influences. Existing studies are confined to quantitative survey-based methods, such as examining the correlation between the presence of eating disorders and the frequency of viewing eating broadcast (Kircaburun et al., 2021c, a; von Ash et al., 2023; Kim et al., 2021; Anari and Eghtesadi, 2023; Bachner-Melman et al., 2018). While these studies are meaningful in establishing links between eating disorders and digital food content, they leave the deeper psychological and behavioral dynamics largely unexplored, creating a significant gap in our understanding. These studies suggested further research into the various aspects of digital food content and media types that potentially promote disordered eating (Kircaburun et al., 2021a; Kim et al., 2021). The goal of our research is to shed light on these intricate dynamics and unravel the profound impact these interactions have on individuals’ attitudes, behaviors, and overall well-being. We aim to provide insights into the mechanisms underpinning the sway of digital food content on this population, paving the way for effective and tailored interventions and support mechanisms on digital food content platforms.
@@ 3. Exploratory Study
Before delving deeply into the investigation of how digital food content affects individuals with EDs, we conducted an IRB-approved exploratory study to broadly assess the extent to which they connect with digital food content in relation to their ED symptoms.

### v4

# Impact Ambivalence: How People with Eating Disorders Get Trapped in the Perpetual Cycle of Digital Food Content Engagement

@@ Abstract.
Digital food content could impact viewers’ dietary health, with individuals with eating disorders being particularly sensitive to it. However, a comprehensive understanding of why and how these individuals interact with such content is lacking. To fill this void, we conducted exploratory (N=23) and in-depth studies (N=22) with individuals with eating disorders to understand their motivations and practices of consuming digital food content. We reveal that participants engaged with digital food content for both disorder‑driven and recovery‑supporting motivations, leading to conflicting outcomes. This impact ambivalence, the coexistence of recovery‑supporting benefits and disorder‑exacerbating risks, sustained a cycle of quitting, prompted by awareness of harm, and returning, motivated by anticipated benefits. We interpret these dynamics within dual systems theory and highlight how recognizing such ambivalence can inform the design of interventions that foster healthier digital food content engagement and mitigate post-engagement harmful effects.
journalyear: 2025journalyear: 2025conference: ; ; copyright: none
@@ 1. Introduction
Digital media’s exponential growth has made food content ubiquitous and diversified. On social media, the popularity of food videos (e.g., eating and cooking broadcasts) and images has significantly increased (Kang et al., 2020; Mejova et al., 2016; De Solier, 2005; Adema, 2000). On TikTok, the hashtag #TikTokFood has been posted over 5M times (TikTok, 2025) and viewed over 43 billion times (Kitchens, 2021), and Instagram has over 313 million posts tagged with #foodporn (Instagram, 2024). The online food market—especially app-based delivery that trades on visually enticing food imagery—has kept surging: delivery accounted for 21% of global consumer food service spending in 2024, up from 9% in 2019 (International, 2025).
@@
A growing body of research has examined the impact of increased exposure to various digital food content on eating habits and dietary health (Petit et al., 2016; Kinard, 2016; Ofli et al., 2017; Vermeir and Roose, 2020; Pope et al., 2015; Choi et al., 2024; Wu et al., 2024; Wang et al., 2024). The trend of watching and posting food images and videos with catchy hashtags has been associated with inadequate eating habits (Yan and Corino, 2019; Qutteina et al., 2019), unhealthy food choices (Peng and JEMMOTT III, 2018; Basso et al., 2018), and overeating (Bodenlos and Wormuth, 2013; Halford et al., 2004) through various visual and social factors (Seal et al., 2022). Browsing and ordering food using online shopping and food delivery services are also known to involve unhealthy food choices (Yeo et al., 2017), unhealthy eating (Eu and Sameeha, 2021), and obesity (Stephens et al., 2020). Moreover, many digital food content features take-away food (Kang et al., 2020), and combining eating content with online food delivery services could lead to unhealthy lifestyles.
@@
Recent studies have highlighted that individuals with eating disorders (EDs) engage with digital food content significantly more than the general population (Bachner-Melman et al., 2018) and are particularly sensitive to such content (Jin, 2018; Legenbauer et al., 2004). When coupled with frequent unconscious exposure to digital food content, it could exacerbate ED symptoms, such as binge eating and obsessive food thoughts (Choi et al., 2024; Kinkel-Ram et al., 2022). Despite these findings, there remains a significant gap in understanding why individuals with EDs engage with digital food content and how such engagement ultimately contributes to their disordered eating in their daily lives. In this regard, we aim to shed light on the experience of people with disordered eating behaviors with digital food content by answering the following question: Why and how do people with EDs interact with digital food content?
@@
To answer this question, we conducted two rounds of studies with individuals with eating disorders: one round of an exploratory study (N=23) and another round of an in-depth study (N=22). The exploratory study assesses how likely individuals with EDs are to be concerned about associations between digital food content and disordered eating, and identifies which digital food content types are associated with disordered eating. Notably, we refrained from bringing up digital food content, allowing participants to broach the subject voluntarily. Sixteen participants out of twenty-three voluntarily linked digital food content to their ED symptoms. Our analysis identified two representative digital food content positively associated with ED symptoms: food media (e.g., images and videos) and food delivery apps. In particular, we found that food media played an ambivalent role; sometimes used for managing eating-related urges, yet at other times acting as a trigger for binge eating.
To further understand the underlying dynamics of this ambivalence, we conducted in-depth interviews focusing on ED individuals’ motivations and practices of consuming digital food content. Going beyond prior research’s emphasis on the detrimental impacts of digital food content on ED populations (Kinkel-Ram et al., 2022; Qutteina et al., 2019; Yan and Corino, 2019), our in-depth study examined how this ambivalence permeates participants’ engagement from motivations that draw them to such content, to the impacts it produces, and the cycles of avoidance and relapse it sustains.
This tension echoes the concept of attitudinal ambivalence, the coexistence of positive and negative evaluations toward the same object or behavior (Thompson et al., 2014; Van Harreveld et al., 2009; Schneider and Schwarz, 2017). In our study, we first observed the coexistence of recovery-supporting and disorder-exacerbating impacts within the same digital food content engagement. We term this coexistence impact ambivalence, extending attitudinal ambivalence to the consequence domain of behavior. Importantly, the recognition of such conflicting impacts shapes individuals’ anticipations of future engagements, which then feed into attitudinal ambivalence: they come to both desire engagement, expecting benefits, and resist it, anticipating harms. Such impact ambivalence not only explains individuals’ recurring cycle of quitting and returning to digital food content, but also surfaces design opportunities often overlooked by binary harm-reduction perspectives. Recognizing this ambivalence enables the identification of moments where supportive and harmful effects intersect, informing interventions that guide users’ decisions before engagement and mitigate adverse impacts afterward.
@@
  • We reveal the perceptions, motivations, impacts, and engagement patterns with various digital food content among individuals with eating disorders. Extending beyond prior literature, which primarily focuses on its detrimental impacts on individuals with EDs, our findings provide deeper insights into why and how these individuals engage with such content, highlighting its overlooked recovery-supporting potential.
  • We articulate the cognitive mechanisms underlying how individuals with EDs interact with digital food content by integrating the concept of impact ambivalence into dual systems theory (Section 6.1). While the double-sided impacts of digital food content in the general population indicate distinct positive (e.g., entertainment) and negative (e.g., triggering overeating) outcomes, impact ambivalence reveals the coexistence of conflicting outcomes, specifically recovery-supporting benefits and ED-exacerbating risks. This perspective offers deeper insights into the unique challenges faced by individuals with EDs.
  • Drawing on a thorough understanding of the cognitive mechanisms behind their content engagement, we propose intervention designs that promote informed and mindful interactions with digital food content and mitigate its effects after engagement.
Our study complements prior research on digital food content by foregrounding the experiences of individuals with eating disorders, providing unique insights into their engagement patterns, and informing the design of digital interventions that better reflect their needs and struggles.
@@ 2.2. An Overarching Presence of Digital Food Content
This omnipresence of digital food content underscores the need to explore its impact on eating behaviors, especially among individuals with eating disorders who are especially susceptible to such content (Jin, 2018; Legenbauer et al., 2004). The subsequent section delves into the intricate dynamics of digital food content’s influence on individuals’ eating behaviors, shedding light on its constructive and detrimental effects.
@@ 2.3. Double-sided Impact of Digital Food Content on Eating Behaviors
Digital food content exerts a complex influence on individuals’ eating behaviors, presenting a dual-sided impact characterized by both constructive and detrimental facets (Strand and Gustafsson, 2020; Wang et al., 2024). While many individuals engage with digital food content for inspiration, recipe ideas, and cultural connections to food (De Solier, 2005), the comprehensive impact of such media on individuals with eating disorders, who are much more sensitive to digital food content than the general population (Schienle et al., 2009; Lee et al., 2017), has received limited attention. The relationship between individuals’ consumption of digital food content, including the engagement with eating broadcasts (Mukbang), and the potential escalation of disordered eating behaviors has garnered recognition in prior research, primarily focusing on the general population (von Ash et al., 2023; Anari and Eghtesadi, 2023). However, these investigations have yielded varying insights and have yet to explicitly target individuals with eating disorders.
@@
Conversely, a growing body of research has highlighted the potentially detrimental impact of digital food content, especially when exacerbating unhealthy eating habits (Pope et al., 2015; Spence et al., 2016). The omnipresence of visually appealing, calorie-laden food imagery in digital content can escalate disordered eating behaviors (Bodenlos and Wormuth, 2013; Ventura et al., 2021). Studies have pointed out that exposure to images and videos of excessive food consumption, particularly those featuring eating challenges or binge eating, can inadvertently trigger or reinforce unhealthy eating patterns (Spence et al., 2019). Such content inadvertently glorifies or normalizes overconsumption (Kircaburun et al., 2021c). Additionally, the visual appeal of enticing food displays and the sensory experience of hearing eating sounds may inadvertently intensify cravings and compulsive eating tendencies (Spence et al., 2016; Anjani et al., 2020). Recent research proposed censoring food visuals to mitigate these adverse effects and demonstrated that preventing unconscious consumption of digital food content can help manage eating disorder conditions in daily life (Choi et al., 2024).
@@
The current research landscape lacks a profound understanding of why and how individuals with eating disorders engage with digital food content—a realm encompassing spontaneous interactions and passive influences. Existing studies are confined to quantitative survey-based methods, such as examining the correlation between the presence of eating disorders and the frequency of viewing eating broadcast (Kircaburun et al., 2021c, a; von Ash et al., 2023; Kim et al., 2021; Anari and Eghtesadi, 2023; Bachner-Melman et al., 2018). While these studies are meaningful in establishing links between eating disorders and digital food content, they leave the deeper psychological and behavioral dynamics largely unexplored, creating a significant gap in our understanding. These studies suggested further research into the various aspects of digital food content and media types that potentially promote disordered eating (Kircaburun et al., 2021a; Kim et al., 2021). The goal of our research is to shed light on these intricate dynamics and unravel the profound impact these interactions have on individuals’ attitudes, behaviors, and overall well-being. We aim to provide insights into the mechanisms underpinning the sway of digital food content on this population, paving the way for effective and tailored interventions and support mechanisms on digital food content platforms.
@@ 3. Exploratory Study
Before delving deeply into the investigation of how digital food content affects individuals with EDs, we conducted an IRB-approved exploratory study to broadly assess the extent to which they connect with digital food content concerning their ED symptoms.